### v4

# On location of maximal gradient of torsion function over some non-symmetric planar domains

Abstract.
We investigate the location of the maximal gradient of the torsion function on some non-symmetric planar domains. First, by establishing uniform estimates for narrow domains, we prove that as a planar domain bounded by two graphs of functions becomes increasingly narrow, the location of the maximal gradient of its torsion function tends toward the endpoint of the longest vertical line segment, with smaller curvature among them. This shows that the Saint-Venant’s conjecture on the location of fail points is valid for asymptotically narrow domains. Second, for triangles, we show that the maximal gradient of the torsion function always occurs on the longest sides, lying between the foot of the altitude and the midpoint of that side. Moreover, via nodal line analysis and the continuity method, we demonstrate that restricted on each side, the critical point of the gradient of the torsion function is unique and non-degenerate. Furthermore, by perturbation and barrier argument, we prove that for a class of nearly equilateral triangles, the critical point is closer to the midpoint than to the foot of the altitude, and the norm of the gradient of the torsion function has a larger value at the midpoint than at the foot of the altitude. Third, using the reflection method, we prove that for a non-concentric annulus, the maximal gradient of torsion always occurs at the point on the inner ring closest to the center of the outer ring.
Research of Qinfeng Li was supported by National Key R&D Program of China (2022YFA1006900) and the National Science Fund of China General Program (No. 12471105). Research of Shuangquan Xie was supported by the Changsha Natural Science Foundation (No. KQ2208006). Research of Ruofei yao was supported by the National Natural Science Fund for Youth Scholars (No. 12001543).
1. Introduction
1.1. Background
This paper deals with one of the most fundamental objects in mathematical physics, namely the torsion function, which is the solution to the following equation:
$\displaystyle\begin{cases}-\Delta u=1&\text{in $\displaystyle\Omega$}\\ u=0&\text{on $\displaystyle\partial\Omega$}.\end{cases}$ (1.1)
In the above, $\Omega$ is often assumed to be a planar Lipschitz domain, and the name “torsion function” came from Saint-Venant’s elasticity theory [32], see also [14] for the physical background behind the notion. Although the equation (1.1) is quite elementary, it is an important research object due to its connections with various mathematical branches and other scientific areas such as mechanical engineering and fluid mechanics, and the amount of literature on this topic is still growing nowadays. Recently, the torsion function has received a lot of attention, not only for its crucial role in geometric functional inequalities and shape optimization (see, for example, the two comprehensive monographs [17] and [18], as well as the very recent paper [1]), but also it often serves as a predictor for the localization of extreme values on Laplacian eigenfunctions, leading to a wealth of inspired results [2, 6, 13, 31, 35, 36] and so on. We also refer to [26] for a survey of qualitative properties of torsion functions on convex domains.
The present paper is motivated by the problem of finding the location of points on $\overline{\Omega}$ at which $|\nabla u_{\Omega}|$ achieves its maximum, where $u_{\Omega}$ is the torsion function associated with $\Omega$. Such a problem traces back to Saint-Venant’s conjecture on fail points. In Saint-Venant’s elasticity theory, $\Omega$ represents the cross section of an elastic cylindrical bar, and $|\nabla u_{\Omega}|$ measures the shear stress. The points at which $|\nabla u_{\Omega}|$ reaches its maximum are called “fail points”, as the bar transitions from elastic state to a plastic state when $|\nabla u_{\Omega}|$ is too large. Since $|\nabla u_{\Omega}|^{2}$ is sub-harmonic, fail points must locate on the boundary of $\Omega$. Saint-Venant then made the following conjecture:
Conjecture 1 (Saint-Venant’s Conjecture).
For a convex planar domain that is symmetric about the two coordinate axes, fail points occur at the contact points of the largest inscribed circle.
Kawohl [25] proved Saint-Venant’s conjecture by imposing the additional assumption that the curvature on $\partial\Omega$ is monotonic in the first quadrant. Moreover, under such conditions, the norm of the gradient of torsion is also monotone in the first quadrant. Later, Ramaswamy [30] and Sweers [37, 38] disproved Saint-Venant’s conjecture by constructing domains symmetric about two axes such that the long axis and the short axis are of the same length while the endpoints have different curvatures. Ramaswamy then proposed a modified conjecture in [30], widely embraced by many mechanical engineers:
Conjecture 2 (Ramaswamy’s Conjecture).
For a planar convex domain that is symmetric about the two coordinate axes, the fail points must occur either at the contact points of the largest inscribed circle, or at points with minimal curvature.
Recently, the first and fourth authors [27] disproved Ramaswamy’s conjecture by providing a precise description of the location of fail points for nearly ball domains. They demonstrate that the location of fail points actually depends on the non-local characteristics of the domain, whereas curvature and contact points of the largest inscribed circle are local features.
Finding the location of fail points also has relevance in addressing thermal problems. In fact, the solution $u$ to (1.1) can be viewed as the steady-state temperature within the thermal body $\Omega$, where the heat source is uniformly distributed in $\Omega$ with a unit value, the heat transfer occurs solely through conduction, and the external temperature remains constant, specifically 00 0. Therefore, the quantity
$\max_{\overline{\Omega}}|\nabla u|=\max_{\partial\Omega}|\nabla u|$
signifies the maximum temperature dispersion across the boundary, and thus the location of fail points pinpoints where this maximum dispersion occurs, which is of interest both theoretically and practically. Moreover, in the insightful article [8], the author introduces a thin thermal insulation problem for the purpose of finding the optimal distribution of insulating material in order to maximize the averaged heat of a given domain, and such a problem recently garnered a lot of renewed interest, as seen in the more recent articles [7], [10], [20] and [21]. In the insulation problem model, when the heat source is uniform, it is shown in [21] that for any $C^{2}$ domain which is not a ball, when the total amount of material is bigger than some threshold, then the optimal insulation necessitates covering the entire boundary. When the material is below the threshold, then it is advantageous not to cover the entire boundary. The above $C^{2}$ regularity assumption is recently dropped by Figalli and Zhang, by proving rigidity of Serrin’s overdetermined system over rough domains, see [12]. Notably, an earlier result in [11] reveals that if the total amount of material goes to 00 0, then the optimal distribution of material actually concentrates on points of maximal dispersion of the solution to (1.1), and thus concentrates on fail points. Hence finding the location of fail points can help design the optimal strategy for distributing insulation material.
For more historical remarks on the location of fail points, we refer to [14], [19], [25] and references therein. We should also remark that it is also an important question concerning how large the maximal gradient of the torsion function can be over convex planar domains with fixed area, due to its wide applications in various contexts. In this direction, we refer to the work of Hoskins and Steinerberger [19], which gives so far the optimal gradient upper bound, to our knowledge.
Although significant progress on the location of fail points has been made since Saint-Venant’s conjecture, much remains unknown. First, previous research has primarily focused on convex domains with one or two axes of symmetry, overlooking domains without symmetry or non-convex domains. More importantly, Saint-Venant’s original conjecture does not incorporate curvature. His rationale may have been that the closer boundary points are to the “center” (which we mean the position where $u_{\Omega}$ attains its maximum value), the larger the norm of the torsion function’s gradient becomes. Even though Saint-Venant’s conjecture has been disproved, the partial validity of the conjecture has not attracted sufficient attention after Kawohl’s work [25], since most works focus on counterexamples. In other words, the spirit behind Saint-Venant’s original conjecture has not been fully explored. Whether or not Saint-Venant’s intuition can be formulated into a rigorous mathematical theorem under some reasonable assumptions without curvature information added remains unknown, and this should be a natural question.
Out of these considerations, we aim to study the location of fail points on three distinct categories of domains: narrow domains, triangles, and non-concentric annuli. More specific motivations related to the above concerns and our main results are described below.
1.2. Our results
Motivated by finding out the conditions for the validity of Saint-Venant’s original conjecture without adding curvature assumptions, we first consider the location of fail points on asymptotically narrow domains. We believe that the asymptotically narrow assumption most rigorously aligns with Saint-Venant’s original spirit. We aim to show that as a domain becomes increasingly narrow, the position of fail points should converge to the location postulated by Saint-Venant’s conjecture. This can be justified once we obtain the following proposition, which provides an asymptotic pointwise formula for the gradient of the torsion function as the domain progressively narrows.
Proposition 1.1.
Let $[a,b]$ be a closed interval, $f_{1}\in C^{4}([a,b])$ be convex and $f_{2}\in C^{4}([a,b])$ be concave satisfying $f_{1}(x)=f_{2}(x)=0$ for $x=a,b$. Let
$\Omega_{\epsilon}:=\{(x,y)\in\mathbb{R}^{2}:x\in[a,b],\leavevmode\nobreak\ \epsilon f_{1}(x)<y<\epsilon f_{2}(x)\}$
and $u_{\epsilon}$ be the torsion function in $\Omega_{\epsilon}$, then we have
$\displaystyle|\nabla u_{\epsilon}(x,\epsilon f_{k}(x))|^{2}=\frac{1}{4}\left(f_{2}(x)-f_{1}(x)\right)^{2}\epsilon^{2}+\epsilon^{4}\lambda_{k,2}(x)+O(\epsilon^{6}),\quad k=1,2,$ (1.2)
where $\lambda_{k,2}(x)$ is a quantity relying on second derivatives of $f_{1}$ and $f_{2}$, and its explicit formula is given in section 3.
The formula (1.2) can be derived via formal Taylor expansion of $\tilde{u}(\epsilon,x,y):=u_{\epsilon}(x,\epsilon y)$ defined on $\Omega_{1}$, once we write it down in the following form:
$\tilde{u}(\epsilon,x,y)=\sum_{k=0}^{4}\tilde{u}_{k}(x,y)\epsilon^{k}+\tilde{R}_{4}(\epsilon,x,y).$
Note that the equation which $\tilde{R}_{4}$ satisfies is not uniformly elliptic as $\epsilon\rightarrow 0$ and thus the rigorous proof of (1.2) requires careful analysis. To overcome the difficulty, instead of establishing the precise order of the $W^{1,\infty}$ norm of the remainder term $\tilde{R}_{4}(\epsilon,\cdot)$ in the above Taylor expansion formula of $\tilde{u}$, we give a sharp estimate of the $W^{1,\infty}$ norm of the remainder term in the expansion of $u_{\epsilon}$, by taking advantage of properties of $P$-function and uniform $L^{\infty}$ estimates for torsion function on narrow domains. Such an argument might be applied to study properties of solutions to some other PDEs over narrow domains. The details of proof can be found in section 3.
Now we let $I$ be the set of points on $x$-axis such that at these points the lengths of vertical cross-sections of $\Omega$ achieve the maximum. Since $\Omega$ is convex, $I$ is either an interval or a single point. Due to Proposition 1.1, we thus immediately have the following.
Theorem 1.2.
Let $(\Omega_{\epsilon})_{0<\epsilon\leq 1}$ be as in Proposition 1.1 and $I\subset(a,b)$ be the interval or point on which $f_{2}-f_{1}$ achieves its maximum. Then the distance between the $x$-coordinates of fail points in $\Omega_{\epsilon}$ and $I$ converges to 00 0 as $\epsilon\rightarrow 0$.
The previous theorem only gives the asymptotic information of $x$-coordinates of fail points, thanks to the first term in the right-hand side of (1.2). In order to further determine the $y$-coordinates of fail points, higher order terms of expansion of $u_{\epsilon}$ are needed, and that is why we need to find out the explicit form of $\lambda_{k,2}$ in (1.2). In fact, concerning the case when $I$ is a single point, further computations yield the following theorem.
[…]
Figure 1.1. Planar domain bounded by two graphs
Theorem 1.3.
Let $(\Omega_{\epsilon})_{0<\epsilon\leq 1}$ be as in Proposition 1.1, and we suppose that $h:=f_{2}-f_{1}$ has only one maximal point $x=z_{0}$. If the curvature of $\partial\Omega_{1}$ at $(z_{0},f_{2}(z_{0}))$ is strictly bigger than that at $(z_{0},f_{1}(z_{0}))$, then there exists $\epsilon_{0}>0$ such that when $0<\epsilon<\epsilon_{0}$, $|\nabla u_{\epsilon}(z_{0},\epsilon f_{2}(z_{0}))|<|\nabla u_{\epsilon}(z_{0},\epsilon f_{1}(z_{0}))|$.
For example, fail points of $\partial\Omega_{\epsilon}$ would be close to the point $P_{\epsilon}$ in Figure 1.1 when $\epsilon$ is small.
Theorem 1.2 and Theorem 1.3 indicate that as a planar domain bounded by two graphs of function becomes increasingly narrow, the location of maximal gradient of its torsion function tends toward the endpoints of the longest vertical line segment, with the point of smaller curvature being the preferred location. More precisely from (1.2), we have shown that the distance from the boundary to the center is the first order term determining the location of fail points, while the curvature information is the second order term. Therefore, our results not only realize the spirit of Saint-Venant, but also confirm the following commonly believed folklore spreading in the circle of mechanical engineers (see [14]) for narrow domains: “the fail point occurs near one of the points of contact of the largest inscribed circle, and of these at that one where the boundary is least convex or most concave.”
In the above, we have considered narrow domains bounded below and above by two smooth functions $f_{1}$ and $f_{2}$. The situation wherein the derivatives $f_{i}^{\prime},\,i=1,2$ diverge at points a and b will also be examined and discussed in Section 3, by truncation argument and comparison with thin rectangles.
Next, we consider the location of fail points on triangular regions, which typically represents a class of domains that may not be symmetric about any axis. The reason for studying fail points on triangles is also motivated by the recent advances on the hot spot conjecture, whose theme is to determine the location of extremal points for the second eigenfunction of the Neumann-Laplacian operator. It is one of the most important locating problems in the field of partial differential equations, and most research works on this conjecture have been centered around planar graphs. Notably, groundbreaking studies on triangles, as seen in [3, 4, 5, 9, 23, 24, 28, 29, 33] and references therein, have provided valuable insights into this conjecture. Drawing inspiration from these elegant findings, we delved into finding the location of fail points on triangles.
First, via the moving plane method or reflection method, we obtain the following result.
Theorem 1.4.
For any triangle $\triangle_{ABC}$, fail points are always located on the longest side $\Gamma$ of $\triangle_{ABC}$. Moreover, these points consistently lie on the line segment between the midpoint of $\Gamma$ and the foot of the altitude from the opposite vertex to $\Gamma$.
A natural question that follows is to count the number of fail points on the boundary of a domain $\Omega$, or more generally, number of critical points of the function:
$\displaystyle\partial\Omega\rightarrow\mathbb{R},$
$\displaystyle x\mapsto|\nabla u|^{2}(x),$
where $u$ is the torsion function on $\Omega$. In this direction and on triangular domains, we obtain the following result, via nodal line analysis, local analysis near vertexes, and the continuity method. Our method is inspired by [9], [22] and [23].
Theorem 1.5.
Let $u$ be the torsion function in a triangle $\Omega=\triangle_{ABC}$. Then the critical point of the function $x\in\partial\Omega\mapsto|\nabla u|^{2}(x)$ on each side of $\triangle_{ABC}$ is unique and nondegenerate. As a result, on each side of the triangle $\triangle_{ABC}$, there is at most one fail point.
Remark.
By the $C^{1}$ regularity of the torsion function over triangles, the gradient vanishes at the vertices. Hence, according to the above theorem, restricted to each side of the triangle, the critical point of the gradient of the torsion function is the unique maximal point.
Combining Theorem 1.4 and Theorem 1.5, we have known that restricted on each side, the maximal point of the norm of the gradient of the torsion function is unique, lying between the foot of the altitude and the midpoint of the side. Now we are curious about the following questions:
Question 1.6.
On each side of a given triangle, which is closer to the maximal point, the foot of the altitude or the midpoint?
Question 1.7.
On each side of a given triangle, at which of the two points (the foot of the altitude or the midpoint of the side) does the magnitude of the norm of the gradient of the torsion function have a larger value?
These are more difficult questions to answer. We start from considering triangles perturbed from an equilateral triangle, and without loss of generality we assume that the equilateral triangle has vertices $A=(-\sqrt{3}/3,0)$, $B=(\sqrt{3}/3,0)$ and $C=(0,1)$. Motivated from the Question 1.6, we have:
Theorem 1.8.
Let $\{\Omega_{t}\}_{t\geq 0}$ be the triangle with vertices $A_{t}=(-\sqrt{3}/3,0)$, $B_{t}=(\sqrt{3}/3+t,0)$ and $C_{t}=(0,1)$. Then when $t$ is close to $0+$, the unique fail point of $\Omega_{t}$ is given by $p_{t}=(x(t),0)$ where
$\displaystyle\frac{7}{24}t+O(t^{2})\leq x(t)<\frac{12}{24}t.$ (1.3)
Remark.
Triangles $\Omega_{t}$ above are generated by stretching the base of the equilateral triangle $\Omega_{0}$, such that $A_{t}B_{t}$ is always the longest side(see Figure 1.2 below).
On the one hand, the foot of the altitude and the midpoint of $A_{t}B_{t}$ are $(0,0)$ and $(t/2,0)$ respectively, and thus for $t>0$ small, $|x(t)-\tfrac{t}{2}|<|x(t)-0|$. Therefore, the theorem above does imply that the fail point is closer to the midpoint of the longest side than the foot of the altitude. On the other hand, after an elementary computation, the contact point of the largest inscribed circle on $A_{t}B_{t}$ has coordinates $(c(t),0)$ with $c(t)=5t/12+O(t^{2})$, and hence such a point also lies in our estimated interval given by (1.3). From this point of view, the original Saint-Venant conjecture also has its rationality on triangular shapes.
[…]
Figure 1.2. Stretched equilateral triangles
We can also answer Question 1.7 for the particular triangle discussed above.
Proposition 1.9.
Let $\{\Omega_{t}\}_{t\geq 0}$ be the triangle with vertices $A_{t}=(-\sqrt{3}/3,0)$, $B_{t}=(\sqrt{3}/3+t,0)$ and $C_{t}=(0,1)$, and let $u(t;x,y)$ be the torsion function over $\Omega_{t}$. Then
$\displaystyle u_{y}\left(t;\frac{t}{2},0\right)-u_{y}(t;0,0)\geq\frac{1}{32}t^{2}+O(t^{3}).$
We consider another perturbation of the equilateral triangle, and also motivated by Question 1.6, we have the following result:
Theorem 1.10.
Let $\triangle_{ABC_{t}}$, $t>0$, be a triangle with vertices $A=(-1/\sqrt{3},0)$, $B=(1/\sqrt{3},0)$ and $C_{t}=(t,1)$. Let the torsion function over $\triangle_{ABC_{t}}$ be denoted by $u(t;x,y)$. Then for $t$ close to $0+$, the critical point of $|\nabla u(t;x,0)|^{2}$ is given by $p_{t}=(x(t),0)$ where
$\displaystyle 0<x(t)\leq\frac{5}{12}t+O(t^{2}).$ (1.4)
Remark.
The triangle above is constructed by moving the top vertex of an equilateral triangle to the right. On the base side, the midpoint is given by $(0,0)$ and the foot of the altitude is given by $(0,t)$, and thus $|x(t)-0|<|x(t)-t|$ for $t>0$ being small. Therefore, the theorem above also shows that restricted to the base side, the critical point of the norm of the gradient of the torsion function is closer to the midpoint.
Similarly motivated by Question 1.7, we have
Proposition 1.11.
Let $\triangle_{ABC_{t}}$, $t>0$, be a triangle with vertices $A=(-1/\sqrt{3},0)$, $B=(1/\sqrt{3},0)$ and $C_{t}=(t,1)$. Let the torsion function over $\triangle_{ABC_{t}}$ be denoted by $u(t;x,y)$. Then,
$\displaystyle u_{y}(t;0,0)-u_{y}(t;t,0)\geq\frac{1}{8}t^{2}+O(t^{3}).$
By Proposition 1.9 and Proposition 1.11, we have shown that for a class of nearly equilateral triangles, the norm of the gradient of the torsion function on the base side has a larger value at the midpoint than at the foot of the altitude. Moreover, due to Theorem 1.8 and Theorem 1.10, the maximum norm is closer to the midpoint than to the foot of the altitude. Even for the special cases, the proofs require a very delicate choice of barrier functions and comparative analysis of the second derivatives, and extending the research method from the perturbation results of an equilateral triangle to those of a general isosceles triangle is difficult. Nevertheless, by numerical observations, we conjecture that the above conclusions hold for generic triangles.
The third category of domains we consider comprises non-concentric annuli, which serve as typical examples of domains containing holes. Although an explicit formula for the torsion function can be derived in the case of a concentric annulus due to its radial symmetry, no such explicit formula exists for a non-concentric annulus. Through the reflection method, we establish the following theorem.
Theorem 1.12.
Suppose that $\Omega\subset\mathbb{R}^{2}$ is a non-concentric annulus, then the fail point in $\overline{\Omega}$ is always located at the point on the inner ring closest to the center of the outer ring.
1.3. Outline of the paper
In section 2, we will mainly prove Proposition 1.1, Theorem 1.2 and Theorem 1.3. In section 3, we will mainly prove Theorem 1.4 and Theorem 1.5. In section 4, we will prove Theorem 1.8, Proposition 1.9, Theorem 1.10 and Proposition 1.11. In section 5, we prove Theorem 1.12.
2. Fail points on narrow domains
In this section, our objective is to derive the asymptotic formula for the squared norm of the gradient of the torsion function on $\partial\Omega_{\epsilon}$ as $\epsilon\rightarrow 0$, where $\Omega_{\epsilon}$ is as in Proposition 1.1.
Before proving Proposition 1.1, we give a useful estimate below for the order of $|u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}$, where $u_{\epsilon}$ is the torsion function in $\Omega_{\epsilon}$.
Proposition 2.1.
Let $\Omega_{\epsilon}$ be parameterized as Proposition 1.1 and $u_{\epsilon}$ be the torsion function in $\Omega_{\epsilon}$. Then we have
$|u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}=O(\epsilon^{2}),$
where we have used the notation $O(\epsilon^{k})$ to represent a function $g$ defined in $\Omega_{\epsilon}$ such that $g/\epsilon^{k}$ is uniformly bounded for any $\epsilon>0$.
Proof.
We denote $H_{-}:=\min_{x\in[a,b]}f_{1}(x),H_{+}=\max_{x\in[a,b]}f_{2}(x)$. We introduce an auxiliary function $v_{\epsilon}(x,y)=\frac{1}{2}(\epsilon H_{+}-y)(y-\epsilon H_{-})$, and then we have
$\displaystyle\begin{cases}-\Delta v_{\epsilon}=1,\quad&\text{in $\displaystyle\Omega_{\epsilon}$}\\ v_{\epsilon}\geq 0&\text{on $\displaystyle\partial\Omega_{\epsilon}$}\end{cases}$ (2.1)
Using the maximum principle for $v_{\epsilon}-u_{\epsilon}$, we have $|u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}\leq|v_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}\leq\frac{1}{8}\epsilon^{2}(H_{+}-H_{-})^{2}$. ∎
As a consequence of Proposition 2.1 and $P$-function argument, see for example [34], we can immediately obtain that $|\nabla u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}$ is of order $O(\epsilon)$. This will be crucial to rigorously prove the error estimates in the equality (1.2) in Proposition 1.1.
Proof of Proposition 1.1.
Let $\tilde{u}(\epsilon,x,y)=u_{\epsilon}(x,\epsilon y)$, and then $\tilde{u}$ satisfies
$\displaystyle\begin{cases}\epsilon^{2}(\tilde{u})_{xx}+(\tilde{u})_{yy}=-\epsilon^{2},\quad&\text{in $\displaystyle\Omega_{1}$}\\ \tilde{u}=0,&\text{on $\displaystyle\partial\Omega_{1}$}.\end{cases}$ (2.2)
We write
$\displaystyle\tilde{u}(\epsilon,x,y)=\sum_{k=0}^{4}\tilde{u}_{k}(x,y)\epsilon^{k}+\tilde{R}_{4}(\epsilon,x,y),$ (2.3)
where $\tilde{u}_{0}=\tilde{u}_{1}=\tilde{u}_{3}=0$,
$\tilde{u}_{2}(x,y)=-\frac{1}{2}(y-f_{1}(x))(y-f_{2}(x)),$
and
$\tilde{u}_{4}(x,y)=-\frac{1}{6}y^{3}a_{1}^{\prime\prime}(x)-\frac{1}{2}y^{2}a_{2}^{\prime\prime}(x)+a_{3}(x)y+a_{4}(x),$
with
$\begin{cases}a_{1}(x)=\tfrac{1}{2}(f_{1}(x)+f_{2}(x)),\\ a_{2}(x)=-\tfrac{1}{2}f_{1}(x)f_{2}(x),\\ a_{3}(x)=\tfrac{1}{6}a_{1}^{\prime\prime}(x)(f_{1}^{2}(x)+f_{1}(x)f_{2}(x)+f_{2}^{2}(x))+\tfrac{1}{2}a_{2}^{\prime\prime}(x)(f_{1}(x)+f_{2}(x)),\\ a_{4}(x)=\tfrac{1}{6}f_{1}^{3}(x)a_{1}^{\prime\prime}(x)+\tfrac{1}{2}a_{2}^{\prime\prime}(x)f_{1}^{2}(x)-a_{3}(x)f_{1}(x).\end{cases}$ (2.4)
[…]
By (2.2) and (2.3), $\tilde{R}_{4}$ satisfies the equation
$\displaystyle\begin{cases}\epsilon^{2}(\tilde{R}_{4})_{xx}+(\tilde{R}_{4})_{yy}=-\epsilon^{6}(\tilde{u}_{4})_{xx},\quad&\text{in $\displaystyle\Omega_{1}$}\\ \tilde{R}_{4}=0,&\text{on $\displaystyle\partial\Omega_{1}$}.\end{cases}$ (2.5)
Hence
$\displaystyle u_{\epsilon}(x,y)=\tilde{u}_{2}(x,\epsilon^{-1}y)\epsilon^{2}+\tilde{u}_{4}(x,\epsilon^{-1}y)\epsilon^{4}+R_{4}(\epsilon,x,y),$ (2.6)
where $R_{4}(\epsilon,x,y)=\tilde{R}_{4}(\epsilon,x,\epsilon^{-1}y)$ is the error term, satisfying
$\displaystyle\begin{cases}\Delta R_{4}(\epsilon,x,y)=-\epsilon^{4}(\tilde{u}_{4})_{xx}(x,\epsilon^{-1}y)=O(\epsilon^{4}),\quad&\text{in $\displaystyle\Omega_{\epsilon}$}\\ R_{4}=0,&\text{on $\displaystyle\partial\Omega_{\epsilon}$}.\end{cases}$ (2.7)
There exists a constant $M>0$, such that $\epsilon^{4}Mu_{\epsilon}\pm R_{4}$ is a superharmonic function. Therefore, by the maximum principle and the boundary Hopf Lemma, we have
$\displaystyle|R_{4}|_{L^{\infty}(\Omega_{\epsilon})}\leq\epsilon^{4}M|u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})},\quad\text{and}\quad|\nabla R_{4}|_{L^{\infty}(\partial\Omega_{\epsilon})}\leq\epsilon^{4}M|\nabla u_{\epsilon}|_{L^{\infty}(\partial\Omega_{\epsilon})}.$
Recall that by Proposition 2.1, we have
$|u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}=O(\epsilon^{2}).$
Using the $C^{1}$ regularity of $u_{\epsilon}$ over $\overline{\Omega}_{\epsilon}$ (see for example [15]) and the classical inequality [34, Eq. 6.12], we obtain the uniform upper bound for $|\nabla u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}$ as follows:
$|\nabla u_{\epsilon}|_{L^{\infty}(\partial\Omega_{\epsilon})}=|\nabla u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}\leq 2|u_{\epsilon}|^{1/2}_{L^{\infty}(\Omega_{\epsilon})}=O(\epsilon).$
Hence, we have
$|R_{4}|_{L^{\infty}(\Omega_{\epsilon)}}=O(\epsilon^{6}),\quad|\nabla R_{4}|_{L^{\infty}(\partial\Omega_{\epsilon})}=O(\epsilon^{5}).$
In view of the expression of $u_{\epsilon}$ given by (2.6) and after some slightly complex but rather straightforward calculations, we have
$|\nabla u_{\epsilon}(x,\epsilon f_{k}(x))|^{2}=\epsilon^{2}\lambda_{k,1}(x)+\epsilon^{4}\lambda_{k,2}(x)+O(\epsilon^{6}),\quad k=1,2,$
where
$\displaystyle\begin{cases}\lambda_{k,1}=(f_{2}-f_{1})^{2}/4\\ \lambda_{k,2}=(a_{1}^{\prime}f_{k}+a_{2}^{\prime})^{2}+(f_{k}-a_{1})(f_{k}^{2}a_{1}^{\prime\prime}+2f_{k}a_{2}^{\prime\prime}-2a_{3})\end{cases}\quad k=1,2.$ (2.8)
Plugging the formulas of $a_{i},i=1,2,3$ given by (2.4) into (2.8), we complete the proof. ∎
Proof of Theorem 1.2.
Using the formula already proved in Proposition 1.1, we have Theorem 1.2. ∎
Proof of Theorem 1.3.
Since $h:=f_{2}-f_{1}$ attains its maximum at $x=z_{0}$, $f_{1}^{\prime}(z_{0})=f_{2}^{\prime}(z_{0})$. Using this fact, we have
$(a_{1}^{\prime}f_{1}+a_{2}^{\prime})^{2}(z_{0})=(a_{1}^{\prime}f_{2}+a_{2}^{\prime})^{2}(z_{0}).$
We let $\Lambda(z_{0})=|\nabla u_{\epsilon}(z_{0},f_{2}(z_{0}))|^{2}-|\nabla u_{\epsilon}(z_{0},f_{1}(z_{0}))|^{2}$. By direct computation, we have
$\displaystyle\Lambda$ $\displaystyle=\epsilon^{4}\{(f_{2}-a_{1})(f_{2}^{2}a_{1}^{\prime\prime}+2f_{2}a_{2}^{\prime\prime}-2a_{3})-(f_{1}-a_{1})(f_{1}^{2}a_{1}^{\prime\prime}+2f_{1}a_{2}^{\prime\prime}-2a_{3})\}+O(\epsilon^{6})$
$\displaystyle=\frac{1}{6}\epsilon^{4}a_{1}^{\prime\prime}(f_{2}-f_{1})^{3}+O(\epsilon^{6})$
$\displaystyle=\frac{1}{12}\epsilon^{4}(f_{1}^{\prime\prime}+f_{2}^{\prime\prime})(f_{2}-f_{1})^{3}+O(\epsilon^{6}).$
Therefore, $|\nabla u_{\epsilon}(z_{0},\epsilon f_{2}(z_{0}))|<|\nabla u_{\epsilon}(z_{0},\epsilon f_{1}(z_{0}))|$ provided that $f_{1}^{\prime\prime}(z_{0})+f_{2}^{\prime\prime}(z_{0})<0$, which is equivalent to $k(z_{0},f_{2}(z_{0}))>k(z_{0},f_{1}(z_{0}))$ where $k$ is the curvature function on $\partial\Omega_{1}$. ∎
When working with convex regions featuring smooth boundaries, we will inevitably face situations where a support line is perpendicular to the $x$-axis. This means that the derivative of $f_{i}$ diverges at the endpoints of $[a,b]$. In this scenario, we have the following two theorems.
Theorem 2.2.
Let $\Omega_{\epsilon}:=\{(x,y)\in\mathbb{R}^{2}:\epsilon f_{1}(x)<y<\epsilon f_{2}(x)\}$ be a family of bounded convex domains with $\partial\Omega\in C^{2}$ and $u_{\epsilon}$ be the torsion function associated with $\Omega_{\epsilon}$. Then for any $\delta>0$, there exists $\epsilon_{0}>0$ such that when $\epsilon<\epsilon_{0}$, we have
$\displaystyle|\nabla u_{\epsilon}(x,\epsilon f_{k}(x))|^{2}=\frac{1}{4}(f_{1}(x)-f_{2}(x))^{2}\epsilon^{2}+O(\epsilon^{3}),\quad\text{for }x\in(a+\delta,b-\delta),k=1,2.$ (2.9)
Theorem 2.3.
Let $(\Omega_{\epsilon})_{0<\epsilon\leq 1}$ be a class of bounded domains which can be parameterized as Proposition 1.1. Then the endpoints $(a,0),(b,0)$ cannot be fail points, provided $\epsilon>0$ is small enough.
Proof of Theorem 2.2.
Let
$D_{\epsilon}:=\{a+\delta/2<x<b-\delta/2\}\cap\Omega_{\epsilon}$
and $\eta\in C_{0}^{\infty}$ be a smooth cut-off function defined by
$\displaystyle\eta(x)=\begin{cases}1,\quad&x\in[a+\delta,b-\delta]\\ 0,&x\leq a+\delta/2\quad\text{and}\quad x\geq b-\delta/2\end{cases}$
with $\eta_{x},\eta_{xx}$ uniformly bounded in $[a,b]$. Let $w_{\epsilon}(x,y):=u_{\epsilon}(x,y)-\hat{\varphi_{\epsilon}}(x,y)$ where $\hat{\varphi}_{\epsilon}(x,y):=\tilde{u}_{2}(x,\epsilon^{-1}y)\epsilon^{2}$ and $\tilde{u}_{2}$ is as in the proof of Proposition 1.1. That is,
$\displaystyle\hat{\varphi_{\epsilon}}(x,y)=-\frac{1}{2}(y-\epsilon f_{1}(x))(y-\epsilon f_{2}(x)).$
We have
$\displaystyle\begin{cases}\Delta(w_{\epsilon}\eta)=\Delta w_{\epsilon}\eta+2\nabla w_{\epsilon}\cdot\nabla\eta+w_{\epsilon}\Delta\eta\quad&\text{in $\displaystyle D_{\epsilon}$}\\ w_{\epsilon}\eta=0&\text{on $\displaystyle\partial D_{\epsilon}$}\end{cases}$ (2.10)
Through direct computation, it is determined that $\Delta w_{\epsilon}=O(\epsilon)$, $|\hat{\varphi}_{\epsilon}|_{L^{\infty}(D_{\epsilon})}=O(\epsilon^{2})$ and $|\nabla\hat{\varphi}_{\epsilon}|_{L^{\infty}(D_{\epsilon})}=O(\epsilon)$. Now we estimate $|(w_{\epsilon})_{x}|_{L^{\infty}(D_{\epsilon})}$ and $|w_{\epsilon}|_{L^{\infty}(D_{\epsilon})}$. Recall that in the proof of Proposition 1.1, we have $|u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}=O(\epsilon^{2})$ and $|\nabla u_{\epsilon}|_{L^{\infty}(\Omega_{\epsilon})}=O(\epsilon)$. Since $D_{\epsilon}\subset\Omega_{\epsilon}$, we have $|(w_{\epsilon})_{x}|_{L^{\infty}(D_{\epsilon})}\leq|(u_{\epsilon})_{x}|_{L^{\infty}(D_{\epsilon})}+|(\hat{\varphi}_{\epsilon})_{x}|_{L^{\infty}(D_{\epsilon})}=O(\epsilon)$ and $|w_{\epsilon}|_{L^{\infty}(D_{\epsilon})}\leq|u_{\epsilon}|_{L^{\infty}(D_{\epsilon})}+|\hat{\varphi}_{\epsilon}|_{L^{\infty}(D_{\epsilon})}=O(\epsilon^{2})$ in $D_{\epsilon}$. Hence from (2.10) we have
$\displaystyle\begin{cases}-\Delta(w_{\epsilon}\eta)=O(\epsilon),\quad&\text{in $\displaystyle D_{\epsilon}$}\\ w_{\epsilon}\eta=0,&\text{on $\displaystyle\partial D_{\epsilon}.$}\end{cases}$ (2.11)
Similar to the derivation of the gradient estimate of $R_{4}$ in the proof of Proposition 1.1, we have
$|\nabla(w_{\epsilon}\eta)|_{L^{\infty}(\partial D_{\epsilon})}=O(\epsilon^{2}).$
Hence
$|\nabla w_{\epsilon}|=O(\epsilon^{2}),\quad\text{uniformly on $\displaystyle\partial\Omega_{\epsilon}\cap\{a+\delta<x<b-\delta\}$}.$
In view that $u_{\epsilon}=\hat{\varphi}_{\epsilon}+w_{\epsilon}$, for $k=1,2$ and $x\in(a+\delta,b-\delta)$, we have
$|\nabla u_{\epsilon}(x,\epsilon f_{k}(x))|^{2}=|\nabla\hat{\varphi}_{\epsilon}|^{2}+O(\epsilon^{3})=\frac{1}{4}(f_{1}(x)-f_{2}(x))^{2}\epsilon^{2}+O(\epsilon^{3}).$
This finishes the proof. ∎
In order to prove Theorem 2.3, we study the asymptotic behavior of norm of the gradient of torsion function on thin rectangles.
Proposition 2.4.
Let $1\displaystyle 0<\epsilon<$ 1, $R_{\epsilon}=[0,1]\times[-\epsilon,\epsilon]$ be a rectangle in $\mathbb{R}^{2}$ and $u_{\epsilon}$ be the torsion function in $R_{\epsilon}$. Then for any point $p$ on the shorter sides of $R_{\epsilon}$, we have
$|\nabla u_{\epsilon}(p)|=O(\epsilon^{2}).$
Proof.
First, by the moving plane method, $|\nabla u_{\epsilon}|(p)\leq|\nabla u_{\epsilon}|(0,0)$ for any $p$ in the shorter sides of the rectangle $R_{\epsilon}$.
By separation of variable argument, the torsion function in rectangle $R_{\epsilon}$ is
$u_{\epsilon}(x,y)=\frac{1}{2}x(1-x)-\frac{2}{\pi^{3}}\sum_{n\geq 1}\frac{1-(-1)^{n}}{n^{3}\cosh(n\pi\epsilon)}\sin(n\pi x)\cosh(n\pi y),$
we have
$f(\epsilon):=|\nabla u_{\epsilon}|(0,0)=|(u_{\epsilon})_{x}|(0,0)=\frac{1}{2}-\frac{4}{\pi^{2}}\sum_{k\geq 0}\frac{1}{(2k+1)^{2}\cosh(n\pi\epsilon)}.$
By direct computation, $f(0)=f^{\prime}(0)=0$. Hence $|\nabla u_{\epsilon}|(0,0)=f(\epsilon)=O(\epsilon^{2})$. We complete the proof. ∎
Proof of Theorem 2.3.
Let
$H=\max\bigg{\{}\max_{z\in[a,b]}|f_{1}(z)|,\leavevmode\nobreak\ \max_{z\in[a,b]}|f_{2}(z)|\bigg{\}}.$
Then the rectangle $R_{\epsilon}:=[a,b]\times[-\epsilon H,\epsilon H]$ contains $\Omega_{\epsilon}$ with $(a,0),(b,0)\in\partial R_{\epsilon}\cap\partial\Omega_{\epsilon}$. Let $u_{\Omega_{\epsilon}}$ be the torsion function in $\Omega_{\epsilon}$ and $v_{\epsilon}$ be the torsion function in $R_{\epsilon}$. By maximum principle, Hopf lemma and Proposition 2.4, we have
$\displaystyle|\nabla u_{\Omega_{\epsilon}}(p)|\leq|\nabla v_{\epsilon}(p)|=O(\epsilon^{2})$
for $p=(a,0)$ or $(b,0)$.
On the other hand, the inradius of $\Omega_{\epsilon}$ is bounded below by $c\epsilon$ for some $c>0$. By the Hopf Lemma and the torsion function’s formula on disks, we derive that the norm of the gradient of the torsion function on $\Omega_{\epsilon}$ at the contact point of some maximal disk contained in $\Omega_{\epsilon}$ is bounded below by some $c\epsilon$, with $c>0$. Consequently, fail points cannot occur at $(a,0)$ or $(b,0)$ when $\epsilon>0$ is small enough. ∎
3. location and uniqueness of fail points on triangles
In this section, we study the location and number of fail points in triangles. Theorem 1.4 applies not only to torsion functions but also to the following semilinear equation:
$\displaystyle\begin{cases}-\Delta u=f(u)&\text{in $\displaystyle\Omega$}\\ u=0&\text{on $\displaystyle\partial\Omega$},\end{cases}$ (3.1)
where $u$ represents a positive solution. For the sake of brevity and clarity, we have chosen to provide proofs specifically tailored to torsion functions.
For two points $A$ and $B$, we use the notation $AB$ to represent the closed line segment that connects the two points. The length of this line segment is denoted by $|AB|$.
First, Theorem 1.4 is a direct consequence of Proposition 3.1 and Proposition 3.2 below.
Proposition 3.1.
Suppose that there exist two sides of $\triangle_{ABC}$ with different lengths, then points on the shorter side can never be fail points of $\triangle_{ABC}$.
Proof.
Let $u$ be the torsion function in $\triangle_{ABC}$. Without loss of generality, we assume that $|AC|<|AB|$. To obtain the desired result, we claim that for any point $p$ on the side $AC$, there exists another point $p^{\prime}$ on the side $AB$ such that the norm of the gradient of $u$ at $p$ (denoted $|\nabla u(p)|$) is less than the norm of the gradient of $u$ at $p^{\prime}$ (denoted $|\nabla u(p^{\prime})|$).
By employing suitable transformations and reflections, we can place the vertex $A$ at the origin, place the vertex $B$ on the positive $x$-axis, and place the vertex $C$ in the upper half plane. Additionally, let $C^{\prime}$ be the point on the line segment $AB$ such that the length of the line segment $AC^{\prime}$ is equal to the length of the line segment $AC$. Furthermore, let $Q$ be a point on the line segment $BC$ where the line $AQ$ bisects the angle $\angle A$. As a result, the triangles $\triangle_{AQC^{\prime}}$ and $\triangle_{AQC}$ are symmetric with respect to the line $AQ$. For example, see Figure 3.1.
[…]
Figure 3.1. Fail point on the longest side
For any $(x,y)\in\Delta_{AQC}$, we let
${w}(x,y)={u}(x,y)-{u}(x^{\prime},y^{\prime}),$
where $(x^{\prime},y^{\prime})$ is the mirror of the point $(x,y)$ with respect to the angular bisector line ${A}{Q}$. That is,
$\displaystyle(x^{\prime},y^{\prime})=(x\cos\alpha+y\sin\alpha,x\sin\alpha-y\cos\alpha),$
where $\alpha$ is the interior angle of triangle $ABC$ at vertex $A$. Observing that
$\displaystyle\begin{cases}\Delta w=0\quad&\text{in $\displaystyle\triangle_{AQC}$},\\ w=0\quad&\text{on $\displaystyle AC\cup AQ$},\\ w<0\quad&\text{on $\displaystyle CQ\setminus\{C,Q\}$},\end{cases}$
the strong maximum implies that ${w}$ takes negative values in the triangle $AQC$. Combining this with the fact that ${w}$ vanishes on the side ${AC}$, we obtain by Hopf lemma that for any $p=(r\cos\alpha,r\sin\alpha)\in AC\setminus\{A,C\}$, the following holds:
$\partial_{\nu_{p}}w(p)=\frac{\partial u}{\partial\nu_{p}}(p)-\frac{\partial u}{\partial\nu_{p^{\prime}}}(p^{\prime})=-|\nabla u|(p)+|\nabla u|(p^{\prime})>0,$ (3.2)
where $p^{\prime}=(r,0)\in AB$ is the reflection of $p$ about $AQ$.
We have shown that for any point $p$ on $AC$, there exists a point $p^{\prime}$ on $AB$ ensuring $|\nabla u(p)|<|\nabla u(p^{\prime})|$. This implies that the fail points cannot exist on the line $AC$. ∎
In the following, when we say critical point of $|\nabla u|^{2}$ on the side $\Gamma$ of a triangle, we mean the critical point of the function
$\displaystyle\Gamma\rightarrow\mathbb{R},$
$\displaystyle x\mapsto|\nabla u|^{2}(x).$
Proposition 3.2.
Let $u$ be the torsion function in the triangle $\triangle_{ABC}$. Then the critical points of $|\nabla u|^{2}$ on each side $\Gamma$ of $\triangle_{ABC}$ exist and must fall along the line segment defined by the midpoint of $\Gamma$ and the foot of the altitude of $\Gamma$.
Proof.
Without loss of generality, we focus on the proof of the theorem on the side $AB$. Since $|\nabla u|=0$ at $A$ and $B$, and $|\nabla u|>0$ in the interior of the side $AB$, the maximum point of $|\nabla u|^{2}$ on the side $AB$ is attained at some interior point of the side $AB$. In particular, there exist at least one critical point of $|\nabla u|^{2}$ on $AB$.
Let $M$ be the midpoint of $AB$ and $F$ be the foot of the altitude on $AB$. Without loss of generality, we assume that $C$ lies in the upper plane, and $A=(0,0)$, $B=(L,0)$ and $F=(l,0)$ where $l\leq L/2$. Let $Q=(L/2,h)$ be the point on the side $BC$ such that $QM\perp AB$. For example, see Figure 3.2.
$A$$B$$C$${F}$${Q}$${M}$
Figure 3.2. Location of fail point
Applying the Hopf lemma on $u$, one obtains that $u_{x}<0$ on $BQ\setminus\{B,Q\}$. Using the same discussion as in the proof of Proposition 3.1, one can deduce that
$u_{x}<0\text{ in $\displaystyle\Delta_{MBQ}$ }$
and $u_{x}\leq 0$ on $QM$, with the equality being valid if and only if $F=M$. Since
$\displaystyle\begin{cases}\Delta u_{x}=0\quad&\text{in $\displaystyle\Delta_{MBQ}$},\\ u_{x}=0\quad&\text{on $\displaystyle MB$},\\ u_{x}<0\quad&\text{inside $\displaystyle\Delta_{ABQ}$},\end{cases}$
by Hopf lemma, $u_{xy}(p)<0$ provided $p\in MB\setminus\{B,M\}$, and hence $|\nabla u|^{2}=u_{y}^{2}$ does not admit any critical point on $MB\setminus\{B,M\}$.
The proof is obviously valid if the interior angle at $A$ is non-acute. Now, we turn to the case where the interior angle at $A$ is acute. By the same reflection argument, in view that the reflection of $\Delta_{AFC}$ about the segment $CF$ is still inside $\Delta_{ABC}$, we have that the harmonic function $u_{x}$ is positive in the triangle $CAF$, and consequently $u_{xy}(p)>0$ for $p\in AF\setminus\{A,F\}$. Therefore, $|\nabla u|^{2}$ does not admit the critical points on $AF\setminus\{A,F\}$, and thus the critical points must lie on $FM$. ∎
Proof of Theorem 1.4.
In view that each fail point in the triangle $\triangle_{ABC}$ must be a critical point of the squared norm of the gradient of torsion function on some side, combining with Proposition 3.1 and Proposition 3.2, we establish the proof. ∎
$A$$B$$C$${C_{\lambda}}$$l:\leavevmode\nobreak\ y=\lambda+x\tan\tfrac{\theta}{2}$
Figure 3.3. The moving plane method for the semilinear case
Remark 3.3.
Proposition 3.1 and Proposition 3.2 hold for the more general semi-linear equation (3.1), via moving plane method instead of reflection. The proof is standard, so we omit it. The Figure 3.3 illustrated the general approach.
Next, we will prove Theorem 1.5 by establishing Lemma 3.4 and Proposition 3.5 below.
Lemma 3.4.
Let $u$ be the torsion function in $\Omega=\triangle_{ABC}$. Suppose that $(u_{\nu})_{\tau}(P)=0$ for some point $P$ in the interior of the side $AB$, where $\tau$ and $\nu$ are the tangential and normal directions corresponding this side $AB$ respectively, then $(u_{\nu})_{\tau\tau}(P)\neq 0$.
Proof.
Without loss of generality, we may suppose that $\triangle_{ABC}\subset\mathbb{R}^{2}_{+}$, where +$\mathbb{R}^{2}_{+}=\{(x,y):y>0\}$ is the upper plane, and we may assume that $AB$ lies in the $x$-axis. Let $\tilde{\Omega}$ be the double domain (kite) of $\Omega$, i.e., the interior of union of $\overline{\Omega}$ and its reflection with respect to the side ${A}{B}$. Let $v$ be the odd extension of $u_{x}$ about $AB$, i.e., $v$ is a function defined in the double domain $\tilde{\Omega}$ as follows:
$\displaystyle v(x,y)=\begin{cases}u_{x}(x,y)&\text{ when }y>0,\\ 0&\text{ when }y=0,\\ -u_{x}(x,-y)&\text{ when }y<0.\end{cases}$
It follows that $v$ is a harmonic function in the double domain $\tilde{\Omega}$ and $v\not\equiv 0$. Since ${v}_{y}({P})={u}_{xy}({P})=0$ and $v_{x}(P)=u_{xx}(P)=0$, $P$ is a zero of $v$ with order at least two. According to [16, Proposition 4.1], $v$ cannot vanish at $P$ with infinite order, and $\{v=0\}$ has at least four branches at $P$ in $\tilde{\Omega}$. Due to the nodal structure of $u_{x}$ at $P$ and its odd symmetry about $AB$, there must exist a nodal line $\mathcal{N}$ of $u_{x}$ inside the triangle ${ABC}$. From the equation of $u$, we have $v_{xx}(P)=0$ and
$v_{yy}(P)=(u_{yy})_{x}(P)=(-1-u_{xx})_{x}(P)=-u_{xxx}(P)=0.$
If the conclusion of the lemma is not true, then $u_{yxx}(P)=0$, and thus
$v_{xy}(P)=u_{xxy}(P)=0.$
Hence all third order derivatives of $v$ vanish, and again by [16, Proposition 4.1], the nodal line of $v$ has at least 66 66 branches at $P$ in $\tilde{\Omega}$, including $PA$ and $PB$, and thus by symmetry, $v$ has at least two nodal lines starting from $P$ going inside $\Delta_{ABC}$.
$A$$B$$C$$P$
Figure 3.4. Nodal line analysis
By Hopf Lemma, $u_{x}\neq 0$ on the interior of the left side $AC$ and the right side $BC$. Consequently, according to the nodal structure of planar harmonic functions, each nodal line of $v$ cannot terminate within $\overline{\Omega}$ before reaching the boundary. Furthermore the hitting point must be one of the vertices. By the strong maximum principle, nodal lines of $v$ cannot form a loop, and hence each nodal line from $P$ going inside $\Delta_{ABC}$ must terminate at the vertex $C$. This again leads to a contradiction, as there are at least two such nodal lines of $v$ that form a loop, with both terminating at $C$. This configuration violates the strong maximum principle, as shown in Figure 3.4. ∎
Consequently, we can establish the uniqueness result as follows.
Proposition 3.5.
Let $u$ be the torsion function in $\triangle_{ABC}$, then the critical point of $|\nabla u|^{2}$ on each side is unique.
Proof.
It suffices to prove the uniqueness of critical point of $|\nabla u|^{2}$ on the side $AB$. After rigid transformation if necessary, we may assume that the side ${A}{B}$ lies on the $x$-axis, and the vertex $C$ lies in the upper plane. Set $A=(-L,0)$, $B=(L,0)$, and $C=(\sigma,h)$, with $h>0$. We may assume that the length of side $AC$ is less than or equals the length of side $BC$, and hence $\sigma\leq 0$.
We first consider the special case $\sigma=0$. In this case, $\triangle_{ABC}$ is an isosceles triangle and the middle point of $AB$ coincides with the foot of the altitude on $AB$. Using Theorem 1.4, the critical point of the gradient of the torsion function on $AB$ is unique, and by Lemma 3.4, it is nondegenerate.
[…]
Figure 3.5. Continuity method via domain deformation
Now we turn to the case $\sigma<0$. We set $C_{t}=(t\sigma,h)$ and denote by $\Omega_{t}$ the triangle $ABC_{t}$; see Figure 3.5. We introduce the following continuous linear transformation:
$F_{t}:(x,y)\mapsto\left(x+t\frac{\sigma}{h}y,y\right),$
which fixes the line segment $AB$ and maps $\Omega_{0}$ to $\Omega_{t}$. Let $u(t;\cdot)$ be the torsion function in $\Omega_{t}$, and let $v(t;x,y)=u(t;F_{t}(x,y))$ which is defined in $\Omega_{0}$. Therefore,
$\displaystyle v_{x}(t;x,y)=u_{x}(t;F_{t}(x,y)),\quad v_{xx}(t;x,y)=u_{xx}(t;F_{t}(x,y)),$
$\displaystyle v_{xy}(t;x,y)=t\frac{\sigma}{h}u_{xx}(t;F_{t}(x,y))+u_{xy}(t;F_{t}(x,y)),$
$\displaystyle v_{y}(t;x,y)=t\frac{\sigma}{h}u_{x}(t;F_{t}(x,y))+u_{y}(t;F_{t}(x,y)),$
and
$v_{yy}(t;x,y)=t^{2}\frac{\sigma^{2}}{h^{2}}u_{xx}(t;F_{t}(x,y))+2t\frac{\sigma}{h}u_{xy}(t;F_{t}(x,y))+u_{yy}(t;F_{t}(x,y)).$
Hence $v$ satisfies
$\displaystyle\begin{cases}\left(1+t^{2}\frac{\sigma^{2}}{h^{2}}\right)v_{xx}+v_{yy}-2t\frac{\sigma}{h}v_{xy}=-1&\text{ in }\Omega_{0},\\ v=0&\text{ on }\partial\Omega_{0}.\end{cases}$
By standard elliptic regularity theory and implicit function theorem, $t\mapsto v(t;\cdot)$ is a smooth map from $[0,1]$ to $C^{1}(\overline{\Omega}_{0})$, and hence $u(t;x,y)$ is smoothly depending on $t$.
In order to show that $|\nabla u|^{2}$ has a unique critical point inside the side $AB$, we consider $F^{t}(x):=|\nabla u(t;x,0)|^{2}$. Recalling that $x=0$ is the unique critical point of the map $F^{0}$ in $(-L,L)$, which is a maximal point, we let $t_{0}$ be the supremum of those values $t\in[0,1]$ such that $F^{t}$ admits a unique critical point in $(-L,L)$. Hence to show that $|\nabla u|^{2}$ has a unique critical point inside $AB$, it suffices to prove that $t_{0}=1$. We prove this by contradiction and suppose
$10\leq t_{0}<$ 1 .
We now do some local analysis for the torsion function near the vertex $A$. We denote the interior angle at vertex $A$ of triangle $ABC_{t}$ by $\alpha_{t}$. We aim to show that near the vertex $A$, there are no critical points of $F^{t}(x)$ when $t$ is close to $t_{0}$.
Case 1: If $\alpha_{t_{0}}<\pi/2$, then there exist $\epsilon_{1}>0$ and $\delta_{1}>0$ such that for any $t$ satisfying $|t-t_{0}|<\epsilon_{1}$, we have $2\displaystyle 0<\alpha_{t}<\pi/$ 2 and $|AH_{t}|>\delta_{1}$, where $H_{t}$ is the point on $AB$ such that $C_{t}H_{t}\perp AB$. By the standard moving plane method, $u_{x}(t;x,y)>0$ for any $(x,y)$ in the interior of $\Delta_{AH_{t}C}$. By Hopf Lemma, particularly we have $u_{xy}(t;x,0)>0$ for any $x\in(-L,\delta_{1}-L]$ and $|t-t_{0}|<\epsilon_{1}$.
Case 2: If $\alpha_{t_{0}}=\pi/2$, then we can take any fixed ${d}\in(0,{h})$, and a point ${R}_{t}=(-L+d\cot\alpha_{t},d)$ in the interior of the left side $AC_{t}$. Since $u_{x}(t_{0};-L,d)>0$ by Hopf lemma, we can choose some small $\delta_{2}>0$ so that the point $(-L+3\delta_{2},d)$ is contained in interior of $\triangle_{ABC}$ and $u_{x}(t_{0};x,d)>0$ for any $x\in[-L,-L+3\delta_{2}]$. Then we can choose $\epsilon_{2}>0$ such that (i) the point $(-L+3\delta_{2},d)$ is still contained in the interior of $\triangle_{ABC_{t}}$ for $|t-t_{0}|<\epsilon_{2}$; (ii) $d|\cot\alpha_{t}|<\delta_{2}$ for $|t-t_{0}|<\epsilon_{2}$; (iii) $u$ satisfies
$u_{x}(t;x,d)>0\text{ for }x\in[-L+d\cot(\alpha_{t}),-L+3\delta_{2}],$
whenever $t\in[t_{0},t_{0}+\epsilon_{2})$. Let ${P}_{\lambda}=(\lambda,d)$ and ${Q}_{\lambda}=(\lambda,0)$, with $\lambda\in(-L,L+\delta_{2}]$. Consequently, the reflection of the open trapezoid $\Omega^{t}_{d,\lambda}$ with vertexes $A,R_{t},{P}_{\lambda},{Q}_{\lambda}$ about the vertical line ${P}_{\lambda}{Q}_{\lambda}$ still lies inside the rectangle $(-L,-L+3\delta_{2})\times(0,d)$. Then for any $\lambda\in[-L,-L+\delta_{2}]$, the function
$w^{\lambda}(t;x,y)=u(t;2\lambda-x,y)-u(t;x,y)$
satisfies
$\displaystyle\begin{cases}\Delta w^{\lambda}(t;\cdot)=0&\text{ in}\Omega^{t}_{d,\lambda},\\ w^{\lambda}(t;\cdot)\geq,\not\equiv 0&\text{ on }\partial\Omega^{t}_{d,\lambda}.\end{cases}$
Hence $w^{\lambda}(t;x,y)>0$ in $\Omega^{t}_{d,\lambda}$. Since $w^{\lambda}(t;\cdot)=0$ on $\{-L+\lambda\}\times[0,d]$, by Hopf Lemma, $w^{\lambda}_{x}(t;\lambda,y)=2u_{x}(t;\lambda,y)>0$ for $y\in(0,d)$. Now that
$\displaystyle\begin{cases}\Delta u_{x}(t;\cdot)=0&\text{ in }\Omega^{t}_{d,\lambda},\\ u_{x}(t;\cdot)>0&\text{ in }\Omega^{t}_{d,\lambda},\\ u_{x}(t;\cdot)=0&\text{ on the side }AB.\end{cases}$
Again by Hopf Lemma, we have $u_{xy}(t;\lambda,0)>0$ for any $t\in[t_{0},t_{0}+\epsilon_{2})$ and $0<\lambda+L\leq\delta_{2}$.
Case 3: If $\alpha_{t_{0}}>\pi/2$, then there exists $\epsilon_{3}>0$ such that when $|t-t_{0}|\leq\epsilon_{3}$, $\alpha_{t}=\angle C_{t}AB\in[\pi/2+\epsilon_{3},\pi-\epsilon_{3}]$. By [15, Chapter 4], (see also the expansion used in [22] and [23] at vertexes), near the vertex $A$ and $|t-t_{0}|\leq\epsilon_{3}$, we can write
$\displaystyle u(t;x,y)={c}_{1}^{t}r^{\omega_{t}}\sin(\omega_{t}\theta)+{c}_{2}^{t}\tilde{x}^{2}+{c}_{3}^{t}\tilde{x}\tilde{y}+{c}_{4}^{t}\tilde{y}^{2}+o(2),$
where $\omega_{t}=\pi/\alpha_{t}$, $r=\sqrt{\tilde{x}^{2}+\tilde{y}^{2}}$, $\tilde{x}=x+L=r\cos\theta$, $\tilde{y}=y=r\sin\theta$ and $o(m)$ represents an infinitesimal of order bigger than $m$ with respect to $r$. The line passing through $A$ and $C_{t}$ is thus given by the equation $y=k(t)(x+L)$, with $k(t)=\tan\alpha_{t}$. Using the equation
$\displaystyle\begin{cases}-\Delta u(t;x,y)=1,\quad(x,y)\in\Delta_{ABC_{t}}\\ u(t;x,k(t)(x+L))=0\quad\text{and}\quad u(t;x,0)=0,\end{cases}$
we have that near $A$,
$\displaystyle u(t;x,y)={c}_{1}^{t}r^{\omega_{t}}\sin(\omega_{t}\theta)+\frac{1}{2}\tilde{y}(k(t)\tilde{x}-\tilde{y})+o(2).$
On the one hand, $\tilde{y}(k(t)\tilde{x}-\tilde{y})<0$ for any point close to $A$ and inside $\triangle_{ABC_{t}}$. On the other hand, $u(t;x,y)$ is a positive function inside $\triangle_{ABC}$. Therefore, it forces that
${c}_{1}^{t}>0\text{ if }|t-t_{0}|\leq\epsilon_{3}.$
By direct computation, there exists some $\rho>0$ such that when $(x,y)\in B_{\rho}(A)$, we have
$\displaystyle u_{yx}(t;x,y)={c}_{1}^{t}\omega_{t}(\omega_{t}-1)r^{\omega_{t}-2}\cos\big{(}(\omega_{t}-2)\theta\big{)}+\frac{k(t)}{2}+o(1).$
Since for any $|t-t_{0}|\leq\epsilon_{3}$, $\omega_{t}\in(1,2)$ and ${c}_{1}^{t}>0$, there exists $\delta_{3}>0$ depending on $\epsilon_{3}$ such that when $-L<x\leq\delta_{3}-L$ and $|t-t_{0}|\leq\epsilon_{3}$,
$\displaystyle u_{yx}(t;x,0)=$ $\displaystyle{c}_{1}^{t}\omega_{t}(\omega_{t}-1)r^{\omega_{t}-2}+\frac{k(t)}{2}+o(1)$
$\displaystyle$>> $\displaystyle{c}_{1}^{t}\omega_{t}(\omega_{t}-1)\delta_{3}^{\omega_{t}-2}+\frac{k(t)}{2}+o(1)>0.$
The above estimate is valid, due to that
$1<\inf_{|t-t_{0}|\leq\epsilon_{3}}\omega_{t}\leq\sup_{|t-t_{0}|\leq\epsilon_{3}}\omega_{t}<2,$
$\inf_{|t-t_{0}|\leq\epsilon_{3}}{c}_{1}^{t}>0,$
and that $o(1)$ has a uniform bound independent of $t$. Actually, by [15, Chapter 4], the remainder term $o(1)$ above only depends on $\|u(t;\cdot)\|_{L^{\infty}(\Omega_{t})}$, which is independent of $t$ since we have the following estimate, due to Talenti [39]:
$\|u(t;\cdot)\|_{L^{\infty}(\Omega_{t})}\leq\|w\|_{L^{\infty}(B)},$
where $B$ is the disk with area $|\Omega_{t}|=|\Omega_{0}|$ and $w$ is the torsion function in the disk $B$.
We can do the same analysis as above for the vertex $B$. In summary, in any case, one can always find $\epsilon_{4}>0$ and $\delta_{4}>0$ such that $|u_{xy}(t;x,0)|>0$ for $x\in(-L,\delta_{4}-L]\cup[L-\delta_{4},L)$ and $|t-t_{0}|<\epsilon_{4}$. In particular, any critical point of $F^{t}$ in $(-L,L)$ must lie in $(-L+\delta_{4},L-\delta_{4})$ when $|t-t_{0}|<\epsilon_{4}$.
Let $x_{t}$ be the unique critical point of $F^{t}$ in $(-L,L)$, for $t<t_{0}$. Then $\partial_{x}F^{t}(x)>0$ for $x\in(-L,x_{t})$ and $\partial_{x}F^{t}(x)<0$ for $x\in(x_{t},L)$. By passing to a subsequence if necessary, we may assume that $x_{t}$ converges to some point $x_{t_{0}}\in[-L,L]$ as $t\uparrow t_{0}$. Hence $x_{t_{0}}$ is a critical (maximal) point of $F^{t_{0}}$, and by the above analysis near vertices, $x_{t_{0}}\in(-L+\delta_{4},L-\delta_{4})$. The nondegeneracy (see Lemma 3.4) of the critical point $x_{t_{0}}$ ensures that $x_{t_{0}}$ is the unique maximal point of $F^{t_{0}}$ and
$\partial_{x}F^{t_{0}}(x)>0\text{ for }x\in(-L,x_{t_{0}})\text{ and }\partial_{x}F^{t_{0}}(x)<0\text{ for }x\in(x_{t_{0}},L).$
The nondegeneracy of $x_{t_{0}}$ also guarantees that there exist two small positive constants $\epsilon_{5}\in(0,\epsilon_{4})$ and $\delta_{5}>0$ such that when $|t-t_{0}|<\epsilon_{5}$, $F^{t}$ has exactly one critical point in $[x_{t_{0}}-\delta_{5},x_{t_{0}}+\delta_{5}]$. By continuity, there exists a small constant $\epsilon_{6}\in(0,\epsilon_{5})$ such that
$\partial_{x}F^{t}(x)>0\text{ for }x\in[-L+\delta_{4},x_{t_{0}}-\delta_{5}]\text{ and }\partial_{x}F^{t}(x)<0\text{ for }x\in[x_{t_{0}}+\delta_{5},L-\delta_{4}]$
for $|t-t_{0}|<\epsilon_{6}$.
Overall, we have shown that when $t\in[t_{0},t_{0}+\epsilon_{6})$, $F^{t}$ has only one critical point in $(-L,L)$, which must lie in $(x_{t_{0}}-\delta_{5},x_{t_{0}}+\delta_{5})$. This yields a contradiction to the definition of $t_{0}$. Hence $t_{0}=1$, and $F^{1}$ has a unique (nondegenerate) critical point. We finish the proof. ∎
Proof of Theorem 1.5.
The result is a combination of Lemma 3.4 and Proposition 3.5. ∎
By Theorem 1.4 and Theorem 1.5, we arrive at the following corollary:
Corollary 3.6.
For any triangle $\triangle_{ABC}$, the number of fail points in $\triangle_{ABC}$ is equal to the number of longest sides. Consequently, $\triangle_{ABC}$ is an equilateral triangle if and only if the number of fail points in $\triangle_{ABC}$ is exactly three.
Last, we remark that through the study of finding the location of fail points on triangles, we obtain another rigidity result for isosceles triangles as a byproduct.
Corollary 3.7.
$\triangle_{ABC}$ is an isosceles triangle if and only if there exists a unit vector $\gamma$ parallel to one side of $\triangle_{ABC}$, such that there is a straight nodal line of $\partial_{\gamma}u$ inside $\triangle ABC$. Here, $u$ denotes the torsion function in $\triangle_{ABC}$.
Proof.
Without loss of generality, we assume that $\gamma$ is parallel to the side $AB$, and $AB$ lies on the $x$-axis. By the nodal structure theorem for planar harmonic functions as used before, the straight nodal line of $u_{x}$ inside $\triangle ABC$ must be the line segment connecting the vertex $C$ and the foot of the altitude on $AB$, denoted by $P$. Using the reflection argument and the Hopf Lemma as before, the line segment $CP$ cannot be the nodal line of $u_{x}$ unless $|AP|=|PB|$. Therefore, $\triangle ABC$ is an isosceles triangle with base $AB$. ∎
4. fail points on nearly equilateral triangles
In this section, we carry on some further analysis on the norm of the gradient of the torsion function on the sides of nearly equilateral triangles. We will prove Theorem 1.8, Proposition 1.9, Theorem 1.10 and Proposition 1.11.
To prove Theorem 1.8, we consider the following smooth mapping given by
$\displaystyle F_{t}(x,y)=\left(\left(1+\frac{\sqrt{3}}{2}t\right)x-\frac{y-1}{2}t,y\right).$ (4.1)
Such $F_{t}$ is constructed to be a linear mapping, fixing $A_{0}$ and $C_{0}$ while moving $B_{0}$ to $B_{t}$. Hence $\Omega_{t}=F_{t}(\Omega_{0})$. By elementary computation, the function $v(t;x,y):=u(t;F_{t}(x,y))$ satisfies
$\displaystyle\begin{cases}\left(1+\frac{t^{2}}{4}\right)\left(1+\frac{\sqrt{3}}{2}t\right)^{-2}v_{xx}+t\left(1+\frac{\sqrt{3}}{2}t\right)^{-1}v_{xy}+v_{yy}=-1,\quad&\textit{in $\displaystyle\Omega_{0}$}\\ v=0,&\textit{on $\displaystyle\partial\Omega_{0}$}\end{cases}$ (4.2)
By standard elliptic regularity theory and the implicit function theorem, $t\mapsto v(t;\textperiodcentered)$ is a smooth map from $[0,1]$ to $C^{1}(\overline{\Omega_{0}})$, and hence $v(t;x,y)$ smoothly depends on $t$. We suppose $v(t;x,y)=v_{0}(x,y)+tv_{1}(x,y)+O(t^{2})$. Then we have
$\displaystyle\begin{cases}-\Delta v_{0}=1,\quad&\textit{in $\displaystyle\Omega_{0}$}\\ v_{0}=0,&\textit{on $\displaystyle\partial\Omega_{0}$}\end{cases}$ (4.3)
and
$\displaystyle\begin{cases}-\Delta v_{1}=(v_{0})_{xy}-\sqrt{3}(v_{0})_{xx},\quad&\textit{in $\displaystyle\Omega_{0}$}\\ v_{1}=0,&\textit{on $\displaystyle\partial\Omega_{0}$}\end{cases}$ (4.4)
Since $v_{0}(x,y)=\frac{y^{3}}{4}-\frac{y^{2}}{2}+\frac{y}{4}-\frac{3x^{2}y}{4}$, we have
$\displaystyle(v_{0})_{xx}(x,y)=-\frac{3y}{2},\quad(v_{0})_{xy}(x,y)=-\frac{3x}{2}.$
Therefore, (4.4) actually is
$\displaystyle\begin{cases}-\Delta v_{1}=\frac{3\sqrt{3}}{2}y-\frac{3x}{2},\quad&\textit{in $\displaystyle\Omega_{0}$}\\ v_{1}=0,&\textit{on $\displaystyle\partial\Omega_{0}$}\end{cases}$ (4.5)
In order to prove Theorem 1.8, we shall derive the formula for the nodal line of $u_{x}$. Using the notation introduced above, we have the following proposition.
Proposition 4.1.
There exists a constant $\delta_{0}>0$ and a function $\varphi(t;y)$, such that the nodal line of $u_{x}(t;\cdot)$ is given by $\mathcal{N}_{t}=\{(\varphi(t;y),y):y\in(0,1)\}$ when $|t|<\delta_{0}$. $\mathcal{N}_{t}$ splits $\Omega_{t}$ into two connected subregions and the contact point $p_{t}$ of $\mathcal{N}_{t}$ with $A_{t}B_{t}$ is the fail point of $\Omega_{t}$. Moreover, we have the formula
$\displaystyle\varphi(t;y)=t\left(\frac{2v_{1}(0,y)}{3y}-\frac{y}{2}+\frac{1}{2}\right)+O(t^{2}),\quad\textit{when $\displaystyle|t|<\delta_{0}$ }$ (4.6)
where $v_{1}$ is the solution to (4.5).
Proof.
By previous analysis in section 3, starting from the fail point $p_{t}\in A_{t}B_{t}$, there is a nodal line $\mathcal{N}_{t}$ of $u_{x}$ going inside the triangle and ending at the vertex $C_{t}$. By the maximal principle, $\mathcal{N}_{t}$ splits $\Omega_{t}$ into two subregions, and on the left of $\mathcal{N}_{t}$, $u_{x}>0$ while on the right of $\mathcal{N}_{t}$, $u_{x}<0$.
Since $v(0;x,y)$ is the torsion function in $\Omega_{0}$, we have $v_{x}(0;0,y)=0,v_{xx}(0;0,y)=-3y/2$. Let
$\displaystyle V(t;x,y):=\begin{cases}\frac{v_{x}(t;x,y)}{y},\quad&y>0\\ v_{xy}(t;x,0),&y=0\\ -\frac{v_{x}(t;x,-y)}{y},&y<0\end{cases}$
Then $V(t;x,y)$ is a smooth function defined in the double domain of $\Omega_{0}$ by reflecting $\Omega_{0}$ about $A_{0}C_{0}$, and $V(0;0,0)=0,V_{x}(0;0,0)=-\frac{3}{2}$. Using the implicit function theorem, there exists a smooth function $\varphi_{0}(t;y)$ with $\varphi_{0}(0;y)=0$ and $\delta^{\prime}_{0}>0$, such that $V(t;x,y)=0$ is uniquely solved by $(\varphi_{0}(t;y),y)$ provided $|t|$ and $|y|$ are less than $\delta^{\prime}_{0}$. By extension and a similar argument, there exists a constant $\delta_{0}>0$ and a smooth function $\psi(t;y),\,y\in(0,1)$ such that $v_{x}(t;\psi(t;y),y)=0,$ for $|t|<\delta_{0}$. Moreover, $\psi(0;y)=0$ and $\psi(t;y)$ depends smoothly on $t$. Since $v_{x}(t;x,y)=(1+\sqrt{3}t/2)u_{x}(t;F_{t}(x,y))$, where $F_{t}(\cdot)$ is defined by (4.1), there exists a function $\varphi(t;y)$ that depends smoothly on $t$, such that
$\mathcal{N}_{t}=\{(\varphi(t;y),y):y\in(0,1)\},\quad\textit{for $\displaystyle|t|<\delta_{0}$}.$
Since $\varphi(t;y)$ depends smoothly on $t$, we may write $\varphi(t;y)=\varphi_{0}(y)+t\varphi_{1}(y)+O(t^{2})$. Noting that $u(0;x,y)$ is symmetric about $y$-axis, we have $\varphi(0;y)=0$, and thus $\varphi_{0}(y)=0$. Using the relation $v_{x}(t;x,y)=(1+\sqrt{3}t/2)u_{x}(t;F_{t}(x,y))$, and since
$\displaystyle F_{t}^{-1}\left(\varphi(t;y),y\right)=$ $\displaystyle\left(\varphi(t;y),y\right)+t\left(\frac{y-1}{2}-\frac{\sqrt{3}}{2}\varphi(t,y),0\right)+O(t^{2})$
$\displaystyle$== $\displaystyle\left(\varphi_{0}(y),y\right)+t\left(\varphi_{1}(y)+\frac{y-1}{2}-\frac{\sqrt{3}}{2}\varphi_{0}(y),0\right)+O(t^{2}),$
we have
00$\displaystyle$ 0 $\displaystyle=v_{x}\left(t;F_{t}^{-1}(\varphi(t;y),y)\right)$
$\displaystyle=(v_{0})_{x}(F_{t}^{-1}(\varphi(t;y),y))+t(v_{1})_{x}(F_{t}^{-1}(\varphi(t;y),y))+O(t^{2})$
$\displaystyle=(v_{0})_{x}(\varphi_{0}(y),y)+t(v_{0})_{xx}(\varphi_{0}(y),y)\left(\varphi_{1}(y)-\frac{\sqrt{3}}{2}\varphi_{0}(y)+\frac{y-1}{2}\right)+t(v_{1})_{x}(\varphi_{0}(y),y)+O(t^{2}).$
Substituting $\varphi_{0}(y)=0$ and the formula of $v_{0}$, we solve for $\varphi_{1}(y)$:
$\displaystyle\varphi_{1}(y)=\frac{2(v_{1})_{x}(0,y)}{3y}-\frac{y}{2}+\frac{1}{2},\quad y\in(0,1)$ (4.7)
This completes the proof. ∎
From Proposition 4.1, the fail point $p_{t}=(x(t),0)$ satisfies
$\displaystyle p_{t}=(\varphi(t;0),0)=(t\varphi_{1}(0)+O(t^{2}),0).$ (4.8)
Here
$\displaystyle\varphi_{1}(0)=\frac{2}{3}(v_{1})_{xy}(0,0)+\frac{1}{2}.$ (4.9)
Now, we are ready to prove Theorem 1.8. The remaining step is to estimate $(v_{1})_{xy}(0,0)$. Since we do not have an explicit formula for $v_{1}$, we have to use the barrier argument. The proof goes as follows.
Proof of Theorem 1.8.
The only remaining proof we need to complete is estimating the value of $(v_{1})_{xy}(0,0)$. Let $w_{1}$ and $w_{2}$ satisfy the following equations, respectively:
$\displaystyle\begin{cases}-\Delta w_{1}=\frac{3\sqrt{3}}{2}y,\quad&\textit{in $\displaystyle\Omega_{0}$}\\ w_{1}=0,&\textit{on $\displaystyle\partial\Omega_{0}$}\end{cases}$ (4.10)
and
$\displaystyle\begin{cases}-\Delta w_{2}=\frac{3}{2}x,\quad&\textit{in $\displaystyle\Omega_{0}$}\\ w_{2}=0,&\textit{on $\displaystyle\partial\Omega_{0}$}\end{cases}$ (4.11)
Then $v_{1}(x,y)=w_{1}(x,y)-w_{2}(x,y)$ and $(v_{1})_{xy}(x,y)=(w_{1})_{xy}(x,y)-(w_{2})_{xy}(x,y)$. We note that $w_{1}(x,y)=w_{1}(-x,y)$, which implies $(w_{1})_{x}(0,y)=0$. Consequently, $(w_{1})_{xy}(0,0)=0$ and thus
$\displaystyle(v_{1})_{xy}(0,0)=(w_{2})_{xy}(0,0).$ (4.12)
Now, we aim to estimate the value of $(w_{2})_{xy}(0,0)$ by comparing the Taylor formula of $w_{2}$ and a suitably chosen barrier function near the origin. For simplicity, we study the behavior of $v_{1}$ in $\triangle_{OB_{0}C_{0}}$, as $w_{2}$ is anti-symmetric with respect to the $y$-axis. We choose the barrier function as follows
$\displaystyle g(x,y)=\frac{1}{8}xy\Big{(}\frac{3}{2}(1-y)^{2}-\frac{9}{2}x^{2}+(1-y)^{3}-3\sqrt{3}x^{3}\Big{)},$ (4.13)
which satisfies
$\displaystyle\begin{cases}-\Delta g=\frac{3}{2}x+\frac{3}{2}xy^{2}+\frac{9\sqrt{3}}{2}x^{2}y,\quad&\textit{in $\displaystyle\triangle_{OB_{0}C_{0}}$}\\ g=0,&\textit{on $\displaystyle\partial\triangle_{OB_{0}C_{0}}$}\end{cases}$
The term $xy$ ensures that $g$ vanishes on the sides $OC_{0}$ and $OB_{0}$, the remaining terms guarantee that $g$ vanishes on the side $B_{0}C_{0}$ as $1-y=\sqrt{3}x$ for $(x,y)$ on $B_{0}C_{0}$.
By the maximum principle, $g(x,y)>w_{2}(x,y)$ for all $(x,y)\in\triangle_{OB_{0}C_{0}}$. Since $w_{2}$ vanishes on the $x$-axis and $y$-axis by boundary condition and anti-symmetry, using the Taylor expansion of $w_{2}$ at the origin, we have
$w_{2}(s,s)=(w_{2})_{xy}(0,0)s^{2}+O(s^{3}),\quad\textit{$\displaystyle 0<s\ll 1$}.$
For $g(x,y)$, we compute:
$g(s,s)=\frac{5}{16}s^{2}+O(s^{3}),\quad\textit{$\displaystyle 0<s\ll 1$}$
Since $g(x,y)>w_{2}(x,y)$ in $\triangle_{OB_{0}C_{0}}$, it follows that
$\displaystyle(w_{2})_{xy}(0,0)\leq\frac{5}{16}.$ (4.14)
Combining (4.8), (4.9), (4.12) and (4.14), we have
$\displaystyle x(t)\geq\left(-\frac{2}{3}\cdot\frac{5}{16}+1/2\right)t+O(t^{2})=\frac{7}{24}t+O(t^{2})$
This completes the proof. ∎
Remark 4.2.
The barrier function given by (4.13) was constructed after extensive trials, and the interval estimation (1.3) of the fail point obtained through this function is currently the most refined version we have achieved.
For a generic triangle $\Omega$ with $u$ being its torsion function, restricted on one side, it is also a natural question to compare the magnitude of $|\nabla u|$ at the foot of the altitude and at the midpoint of the side. For the particular situation considered above, we have
Proposition 4.3.
Let $\{\Omega_{t}\}_{t\geq 0}$ be the triangle with vertices $A_{t}=(-\sqrt{3}/3,0)$, $B_{t}=(\sqrt{3}/3+t,0)$ and $C_{t}=(0,1)$, and let $u(t;x,y)$ be the torsion function over $\Omega_{t}$. Then there exists $\delta>0$ such that when $0<t<\delta$, $u_{y}(t;t/2,0)>u_{y}(t;0,0)$.
It suffices to prove Proposition 1.9.
Proof of Proposition 1.9.
We adopt the notations used in the proof of Theorem 1.8. By direct computation, we have
$\displaystyle F_{t}^{-1}(x,y)=\left(x+t\left(\frac{y-1}{2}-\frac{\sqrt{3}}{2}x\right)+t^{2}\left(\frac{3}{4}x-\frac{\sqrt{3}}{4}(y-1)\right)+O(t^{3}),y\right).$
Since
$\displaystyle u(t;x,y)=v(t;F_{t}^{-1}(x,y)),$
we have
$\displaystyle u_{y}(t;x,y)=v_{x}\left(t;F_{t}^{-1}(x,y)\right)\left(\frac{t}{2}-\frac{\sqrt{3}}{4}t^{2}\right)+v_{y}\left(t;F_{t}^{-1}(x,y)\right)+O(t^{3}).$
Recalling that $v(t;x,y)=v_{0}(x,y)+tv_{1}(x,y)+t^{2}v_{2}(x,y)+O(t^{3})$, the explicit formula for $v_{0}$, Taylor expansion and straightforward computations yield
$\displaystyle u_{y}(t;t/2,0)-u_{y}(t;0,0)=\frac{t^{2}}{2}\left((v_{1})_{xy}(0,0)+\frac{3}{8}\right)+O(t^{3}).$
By the key estimate (4.14), we have
$\displaystyle(v_{1})_{xy}=-(w_{2})_{xy}(0,0)\geq-\frac{5}{16}.$
Therefore,
$\displaystyle u_{y}(t;t/2,0)-u_{y}(t;0,0)\geq\frac{1}{32}t^{2}+O(t^{3}).$
This completes the proof. ∎
Next, we prove Theorem 1.10, and we also let $\Omega_{t}$ be the triangles in the hypothesis and $\mathcal{N}_{t}$ be the nodal line of $u_{x}(t;x,y)$. Similarly as above, we know that for $t>0$ small, $\mathcal{N}_{t}$ can be written as a graph of function of the $y$-variable, also denoted by $\varphi$. We also let $v(t;x,y)=u(t;F_{t}(x,y))$ where $F_{t}(x,y)=(x+ty,y)$. Such $F_{t}$ maps $\Omega_{0}$ to $\Omega_{t}$. We can check that $v(t;\cdot)$ satisfies
$\displaystyle\begin{cases}(1+t^{2})v_{xx}+v_{yy}-2tv_{xy}=-1&\text{ in }\triangle ABC_{0},\\ v=0&\text{ on }\partial\triangle ABC_{0}.\end{cases}$ (4.15)
By standard elliptic regularity theory and implicit function theorem, $t\mapsto v(t;\cdot)$ is a smooth map from $[0,1]$ to $C^{1}(\overline{\triangle ABC_{0}})$, and hence $v(t;x,y)$ is smoothly depending on $t$. We suppose $v(t;x,y)=v_{0}(x,y)+tv_{1}(x,y)+O(t^{2})$, as $t\to 0$. Then we have
$\displaystyle\begin{cases}-\Delta v_{0}=1,\quad&\textit{in }\Omega_{0}\\ v_{0}=0,&\textit{on }\partial\Omega_{0}\end{cases}$
and
$\displaystyle\begin{cases}\Delta v_{1}=2(v_{0})_{xy},\quad&\textit{in }\Omega_{0}\\ v_{1}=0&\textit{on }\partial\Omega_{0}\end{cases}$
Since
$v_{0}(x,y)=\frac{1}{4}y^{3}-\frac{1}{2}y^{2}+\frac{1}{4}y-\frac{3}{4}x^{2}y,$
we have $(v_{0})_{y}(x,0)=\frac{1}{4}-\frac{3}{4}x^{2},(v_{0})_{xy}(x,0)=-\frac{3}{2}x$, and hence
$\displaystyle\begin{cases}-\Delta v_{1}=3x,\quad&\textit{in }\Omega_{0}\\ v_{1}=0,&\textit{on }\partial\Omega_{0}\end{cases}$
With these preparations, the proof of Theorem 1.10 goes as follows.
Proof of Theorem 1.10.
Since $u_{x}(t;\varphi(t;y),y)=0$,
$0=v_{x}(t;\varphi(t;y)-ty,y).$
By symmetry, $\phi(0;y)=0$, and thus we may write
$\displaystyle\varphi(t;y)=t\varphi_{1}(y)+O(t^{2}).$ (4.16)
Then
$0=(v_{0})_{x}(0,y)+t\Big{(}(v_{0})_{xx}(0,y)(\varphi_{1}(y)-y)+(v_{1})_{x}(0,y)\Big{)}+O(t^{2}).$
Since $(v_{0})_{x}(0,y)=0,(v_{0})_{xx}(0,y)=-\frac{3}{2}y$, we have
$0=t\Big{(}-\frac{3}{2}y(\varphi_{1}(y)-y)+(v_{1})_{x}(0,y)\Big{)}+O(t^{2}).$
This gives $\varphi_{1}(y)=\frac{2}{3}\frac{(v_{1})_{x}(0,y)}{y}+y.$ Therefore,
$\displaystyle\varphi_{1}(0)=\frac{2}{3}(v_{1})_{xy}(0,0)$ (4.17)
Now, we estimate the value of $(v_{1})_{xy}(0,0)$. From the equation of $v_{1}$, $v_{1}(x,y)=-v_{1}(-x,y)$, and this implies $v_{1}(0,y)=0$ and $v_{1}>0$ in $\triangle_{OBC_{0}}$.
We let
$w=\frac{1}{4}xy\Big{(}\frac{3}{2}(1-y)^{2}-\frac{9}{2}x^{2}+(1-y)^{3}-3\sqrt{3}x^{3}\Big{)},$
which clearly satisfies
$\displaystyle\begin{cases}-\Delta w=3x+3xy^{2}+9\sqrt{3}x^{2}y,\quad&\textit{in $\displaystyle\triangle_{OBC_{0}}$}\\ w=0,&\textit{on $\displaystyle\partial\triangle_{OBC_{0}}$}\end{cases}$
Since
$\displaystyle\begin{cases}-\Delta v_{1}=3x<\Delta w,\quad&\textit{in $\displaystyle\triangle_{OBC_{0}}$}\\ v_{1}=0,&\textit{on $\displaystyle\partial\triangle_{OBC_{0}}$}\end{cases}$
We have $w(x,y)>v(x,y)$ for all $(x,y)\in\triangle_{OBC_{0}}$. Since $v_{1}$ vanishes on the $x$-axis and $y$-axis and $\Delta v_{1}$ vanishes at the origin, using the Taylor expansion of $v_{1}$ at origin, we have
$v_{1}(s,s)=(v_{1})_{xy}(0,0)s^{2}+O(s^{3}),\quad\textit{$\displaystyle 0<s\ll 1$}.$
For $w$, we have
$w(s,s)=\frac{5}{8}s^{2}+O(s^{3}),\quad\textit{$\displaystyle 0<s\ll 1$}$
Then the fact that $w(x,y)>v_{1}(x,y)$ in $\triangle_{OBC_{0}}$ ensures that
$\displaystyle(v_{1})_{xy}(0,0)\leq\frac{5}{8}.$ (4.18)
Combining (4.16), (4.17) and (4.18), we have
$\varphi(t;0)\leq\frac{5}{12}t+O(t^{2}).$
This completes the proof. ∎
Also, motivated by Question 1.7, we compare the values of $|\nabla u(t;x,y)|$ at the midpoint $(0,0)$ and the foot of the altitude $(t,0)$. We have
Proposition 4.4.
Let $\triangle_{ABC_{t}}$, $t>0$, be a triangle with vertices $A=(-1/\sqrt{3},0)$, $B=(1/\sqrt{3},0)$ and $C_{t}=(t,1)$. Let the torsion function over $\triangle_{ABC_{t}}$ be denoted by $u(t;x,y)$. Then, there exists $\delta>0$ such that when $0<t<\delta$, $u_{y}(t;0,0)>u_{y}(t;t,0)$.
It suffices to prove Proposition 1.11.
Proof of Proposition 1.11.
We adopt the notations used in the proof of Theorem 1.10.
Since
$\displaystyle u(t;x+ty,y)=v(t;x,y),$
we have
$\displaystyle u_{y}(t;x,y)=-tv_{x}(t;x-ty,y)+v_{y}(t;x-ty,y).$
Recalling that $v(t;x,y)=v_{0}(x,y)+tv_{1}(x,y)+t^{2}v_{2}(x,y)+O(t^{3})$, by the expression of $v_{0}$ and Taylor expansion, straightforward computations yield
$\displaystyle u_{y}(t;0,0)-u_{y}(t;t,0)=\left(\frac{3}{4}-(v_{1})_{xy}(0,0)\right)t^{2}+O(t^{3})$
By the key estimate (4.18), we immediately have
$\displaystyle u_{y}(t;0,0)-u_{y}(t;t,0)\geq\left(\frac{3}{4}-\frac{5}{8}\right)t^{2}+O(t^{3})=\frac{1}{8}t^{2}+O(t^{3}).$
This completes the proof. ∎
We should mention that a similar argument as in the proof of Theorem 1.8 and Theorem 1.10 can also derive the location of the maximal point of $|\nabla u(t;x,y)|^{2}$ on other sides of the two types of nearly equilateral triangles considered above, and can also be extended to other perturbed equilateral triangles. Similarly, Proposition 1.9 and Proposition 1.11 may also be extended to other sides of the particular nearly equilateral triangles considered in this section and other nearly equilateral triangles. For brevity but without loss of generality, we only presented these results at present. Question 1.6 and Question 1.7 remain open for generic triangles, and answering them requires new insight.
5. Fail points on non-concentric annuli
In this section, we consider a class of multi-connected domains, annuli. Concentric annuli have symmetric, uniform shapes with equidistant points from a shared center, facilitating straightforward mathematical analysis. In contrast, non-concentric annuli lack this symmetry, resulting in more varied and complex shapes due to differing radial distances from distinct centers. In this section, we study the location of fail points in annuli by the reflection method .
Proof of Theorem 1.12.
Without loss of generality, we assume that the non-concentric annulus $\Omega$ is given by
$\Omega=\{x\in\mathbb{R}^{2}:|x|<\rho_{1},\leavevmode\nobreak\ \leavevmode\nobreak\ |x-{I}|>\rho_{2}\},$ (5.1)
where ${I}=(\epsilon,0)$ and $\rho_{1}$, $\rho_{2}$ and $\epsilon$ are positive constants satisfying $\rho_{2}+\epsilon<\rho_{1}$.
Step 1. First, we can show that fail points must be located on the inner boundary, by proving that for every outer boundary point $p$, there exists a corresponding inner boundary point $q$ such that $|\nabla u({p})|<|\nabla u({q})|$. In fact, we fix any outer boundary point ${p}^{\theta}=(\rho_{1}\cos\theta,\rho_{1}\sin\theta)$. Set ${V}^{\theta}=(\rho\cos\theta,\rho\sin\theta)$ with $\rho=\rho_{1}+\rho_{2}$. Then ${p}^{\theta}$ is the unique point of $\partial{B}_{\rho_{1}}\cap\partial{B}_{\rho_{2}}({V}^{\theta})$. We denote by ${T}^{\theta}$ the perpendicular bisector of the line segment connecting ${V}^{\theta}$ and the center $(\epsilon,0)$ of the inner ball, i.e.,
${T}^{\theta}=\{x\in\mathbb{R}^{2}:2x_{1}(\rho\cos\theta-\epsilon)+2x_{2}\rho\sin\theta=\rho^{2}-\epsilon^{2}\}.$ (5.2)
[…]
Figure 5.1. Non-concentric annulus: inner and outer boundaries
The mirror of the point ${p}^{\theta}$ with respect to ${T}_{\theta}$ is denoted by ${q}^{\theta}=({q}_{1}^{\theta},{q}_{2}^{\theta})$ with
$\displaystyle{q}_{1}^{\theta}$ $\displaystyle=\frac{(\rho^{2}-\epsilon^{2})(\rho\cos\theta-\epsilon)-\rho_{1}(\rho^{2}\cos\theta-2\rho\epsilon+\epsilon^{2}\cos\theta)}{\rho^{2}+\epsilon^{2}-2\rho\epsilon\cos\theta},$
$\displaystyle{q}_{2}^{\theta}$ $\displaystyle=\frac{(\rho^{2}-\epsilon^{2})(\rho-\rho_{1})\sin\theta}{\rho^{2}+\epsilon^{2}-2\rho\epsilon\cos\theta}.$
By the construction, ${q}^{\theta}$ is on the inner boundary. It is clear that
$u(x)-u^{\theta}(x)<0\text{ for }x\in{D}^{\theta},$ (5.3)
where $u^{\theta}(x)=u({T}^{\theta}x)$, ${T}^{\theta}x$ stands for the mirror of $x$ w.r.t. ${T}_{\theta}$, and ${D}^{\theta}$ is the smaller cap cut by ${T}^{\theta}$ from ${B}_{\rho_{1}}$. The Hopf lemma implies that $x\cdot\nabla(u-u^{\theta})>0$ at $x={p}^{\theta}$. It follows that $|\nabla u({p}^{\theta})|<|\nabla u({q}^{\theta})|$.
Step 2. Second, we decide the exact location of the fail point on the inner boundary. Let ${L}_{\varphi}$ be the straight line that passes through the center of the inner circle and has the slant angle $\varphi\in(0,\pi)$. That is ${L}_{\varphi}=\{x\in\Omega:(x_{1}-\epsilon)\sin\varphi+x_{2}\cos\varphi=0\}$. Let
$\tilde{D}_{\varphi}=\{x\in\Omega:(x_{1}-\epsilon)\sin\varphi+x_{2}\cos\varphi>0\}.$ (5.4)
${L}_{\varphi}$$x$$y$${I}$${O}$${q_{0}}$${q_{\varphi}}$
Figure 5.2. Non-concentric annulus: location of fail point
Again one can show that
$u(x)-u_{\varphi}(x)<0\text{ for }x\in\tilde{D}_{\varphi},$ (5.5)
where $u_{\varphi}(x)=u({L}_{\varphi}x)$, ${L}_{\varphi}x$ stands for the mirror of $x$ w.r.t. ${L}_{\varphi}$, and $\tilde{D}_{\varphi}$ is the smaller region cut by ${L}_{\varphi}$ from $\Omega$. The Hopf lemma implies that $(x-{I})\cdot\nabla(u-u_{\varphi})<0$ on $(\partial\tilde{D}_{\varphi}\setminus{L}_{\varphi})\cap\partial{B}_{\rho_{2}}({I})$. In particular,
$|\nabla u({q}_{0})|>|\nabla u({q}_{\varphi})|\text{ for }\varphi\in(0,\pi),$
where ${q}_{\varphi}=(\epsilon-\rho_{2}\cos(2\varphi),-\rho_{2}\sin(2\varphi))$. Hence we complete the proof of Theorem 1.12. ∎
As a consequence of the theorem above, we have:
Remark 5.1.
An annulus has a unique fail point if and only if it is non-concentric.
Remark 5.2.
Using the moving plane method instead of the reflection argument, the result of Theorem 1.12 can be extended to higher dimensions and the semilinear equation (3.1), rather than being limited to torsion functions in two dimensions.